# Redundant Queries in DETR-Based 3D Detection Methods: Unnecessary and Prunable

### v1

@@ Abstract
Query-based models are extensively used in 3D object detection tasks, with a wide range of pre-trained checkpoints readily available online. However, despite their popularity, these models often require an excessive number of object queries, far surpassing the actual number of objects to detect. The redundant queries result in unnecessary computational and memory costs. In this paper, we find that not all queries contribute equally – a significant portion of queries have a much smaller impact compared to others. Based on this observation, we propose an embarrassingly simple approach called Gradually Pruning Queries (GPQ), which prunes queries incrementally based on their classification scores. A key advantage of GPQ is that it requires no additional learnable parameters. It is straightforward to implement in any query-based method, as it can be seamlessly integrated as a fine-tuning step using an existing checkpoint after training. With GPQ, users can easily generate multiple models with fewer queries, starting from a checkpoint with an excessive number of queries. Experiments on various advanced 3D detectors show that GPQ effectively reduces redundant queries while maintaining performance. Using our method, model inference on desktop GPUs can be accelerated by up to 1.31x. Moreover, after deployment on edge devices, it achieves up to a 67.86% reduction in FLOPs and a 76.38% decrease in inference time. The code will be available at https://github.com/iseri27/Gpq.
@@ 1 Introduction
Figure 2: Selection frequency of queries (sorted in ascending order) during inference for different methods. Among these methods, PETR, PETRv2, FocalPETR and StreamPETR are 3D object detection methods, and the other two are 2D object detection methods. As illustrated, the selection frequency of queries is imbalanced across both 2D and 3D query-based methods. In PETR, PETRv2, and FocalPETR, there are even queries that were never selected as final results.
@@ 1.2 Pruning Redundant Queries
    We conducted extensive experiments on various query-based detectors to evaluate the effectiveness of our proposed method. The results indicate that, on desktop GPUs, GPQ achieves inference acceleration of up to 1.31× and it achieves at most a 67.87% reduction in FLOPs with a 76.38% decrease in inference time after deployment on edge devices.
@@ 3.1 Revisiting Query-based Methods
[…]
@@
[…]
@@ 4.1 Experimental Setup
4.1.1 Dataset and Detectors
We conduct our experiments on the nuScenes dataset [2], which consists of over 23,000 samples. Each sample includes images from six surrounding cameras, covering the full 360 field of view. The 3D detection task involves 10 object classes (car, truck, bus, trailer, construction vehicle, pedestrian, motorcycle, bicycle, barrier, and traffic cone.), including both static and dynamic objects.
To validate the effectiveness and efficiency of our proposed method, we perform experiments on five advanced detectors: DETR [48], PETR [32], PETRv2 [33], FocalPETR [46] and StreamPETR [47]. Notably, PETRv2 employs VovNet [22] as its image backbone, while the others utilize ResNet50 [15] as their image backbone.
4.1.2 Evaluation Metrics
In the field of 3D object detection, the primary performance evaluation metrics are mean Average Precision (mAP). NuScenes also provides the nuScenes Detection Score (NDS), which is derived from mAP, along with several other error metrics: mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), and mean Average Attribute Error (mAAE). Mathematically, $\text{NDS}=\cfrac{1}{5}\left[5\text{mAP}+\sum\limits_{\text{mTP}}\left(1-\min(1,\text{mTP})\right)\right]$, where $\text{mTP}\in\{\text{mATE},\text{mASE},\text{mAOE},\text{mAVE},\text{mAAE}\}$.
4.1.3 Implementation Details
We use AdamW [35] optimizer with a weight decay of $1.0\times 10^{-2}$. When training from scratch, we use a base learning rate of $2\times 10^{-4}$ with a batch size of $8$, and apply a cosine annealing policy [36] for learning rate decay. When fine-tuning from an existing checkpoint, the learning rate is set to $1.0\times 10^{-4}$. The point cloud range is configured to $[-61.2\mathrm{m},61.2\mathrm{m}]$ along the $X$ and $Y$ axis, and $[-10.0\mathrm{m},10.0\mathrm{m}]$ along the $Z$ axis for all methods.
@@ 4.1.3 Implementation Details
Table 2: Pruning results for different models. The column of “Queries” is of format “# Initial queries/ # Final queries”. Lines where “# Final queries” remain blank are baselines used to compare. All baselines and checkpoints are trained for 24 epochs, with the pruning process completed within the first 6 epochs when loading a checkpoint. The FPS is measured on a single RTX3090 GPU.
@@ 4.2 Main Results
4.2.1 Detection Accuracy: Before vs. After Pruning
@@
One of the key objectives of our pruning method is to maintain the performance of the original models. As demonstrated in Table 2, our approach successfully preserves, and in some cases enhances, the performance of various detectors, even when pruning a checkpoint with a large number of queries. GPQ can also accelerate model inference on desktop GPUs. For example, in the case of PETR, pruning from 900 to 150 queries results in an mAP of 30.52%, which is 2.15 points higher than training from scratch with 150 queries (28.37%), and its speed increases from 6.9 fps to 9.3, which is 1.35x faster. Remarkably, pruning from 900 to 300 queries achieves an mAP of 32.85%, outperforming the result of training from scratch with 900 queries (31.74%). Similarly, for both FocalPETR and StreamPETR, pruning from 900 to 300 queries yields optimal performance, surpassing the results obtained from training with 900 queries from scratch. For PETRv2, while pruning from 900 to 300 queries results in a slightly lower mAP than training from scratch with 900 queries, it still exceeds the performance of training from scratch with 300 queries.
@@
Inference speed is crucial for deploying models on edge devices. To verify whether pruning queries can indeed improve speed, we export the model to ONNX format and deploy it on the Jetson Nano B01 using ONNX Runtime to measure the model’s running time. After pruning from 900 to 300 queries, the FLOPs of StreamPETR are reduced by 28%, and the running time decreases by 39.74%. Pruning further from 900 to 150 queries results in a 47.96% reduction in running time, making the model 1.92×1.92$\times$1.92 × faster. The pruning method also proves effective for FocalPETR and PETR, significantly reducing both FLOPs and running time. Further details can be found in Table 3.
Since our method does not modify the image backbones or necks, we remove the backbone module to better illustrate the effects of our approach. Specifically, we run only the transformer decoder using randomly generated dummy inputs. Results are also shown in Table 3. For all the evaluated models, pruning from 900 queries to 300 results in saving more than half of the running time for transformer-related modules. Additionally, compared to the model with 900 queries, StreamPETR with 150 pruned queries achieves a 76.38% reduction in running time.
4.2.3 Comparison with ToMe
To the best of our knowledge, this is the first study to explore query redundancy in Transformer-based 3D detection models. As a result, it is challenging to find comparable methods for direct evaluation. However, in this section, we compare our approach with ToMe [1], a method that shares a similar goal of improving efficiency. Specifically, ToMe increases the throughput of Vision Transformer (ViT) models by dividing $N$ tokens into two equal size groups $\mathbb{A}$ and $\mathbb{B}$. These tokens represent image features extracted by the image backbone. To merge similar tokens, ToMe calculates a similarity matrix $\mathbf{S}$ of size $\frac{N}{2}\times\frac{N}{2}$, where the $(i,j)$-th entry represents the similarity between token $i$ in group $\mathbb{A}$ and token $j$ in group $\mathbb{B}$. It then merges $r$ token pairs with the highest similarity scores.
@@ 4.2.3 Comparison with ToMe
Model r Queries Speed(fps) mAP $\uparrow$ NDS $\uparrow$ mATE $\downarrow$ mASE $\downarrow$ mAOE $\downarrow$ mAVE $\downarrow$ mAAE $\downarrow$
StreamPETR - 900/- 16.1 37.83% 0.4734 0.6961 0.2822 0.6846 0.2856 0.2084
StreamPETR-ToMe 264 900/- 14.8 37.79% 0.4731 0.6982 0.2819 0.6849 0.2849 0.2084
StreamPETR-ToMe 528 900/- 14.9 37.69% 0.4721 0.6994 0.2822 0.6855 0.2877 0.2088
StreamPETR-ToMe 1056 900/- 15.4 36.34% 0.4608 0.7178 0.2852 0.6963 0.2965 0.2121
StreamPETR-ToMe 2112 900/- 16.0 31.69% 0.4325 0.7546 0.2907 0.6893 0.3170 0.2210
StreamPETR-GPQ - 900/300 18.7 39.42% 0.4941 0.6766 0.2711 0.5799 0.2780 0.2136
@@
Table 4: Comparison with ToMe. Results indicate that ToMe performs poorly on 3D object detection, while our GPQ effectively maintains the model’s performance and accelerates inference.
@@
We apply ToMe to StreamPETR, and compare its performance with GPQ in Table 4. Surprisingly, instead of accelerating computation, ToMe causes the model to run slower. We believe this is due to the overhead introduced by calculating the similarity matrix. Unlike image classification tasks, which typically handle only a few hundred tokens, 3D object detection methods often generate a much larger number of tokens (e.g., StreamPETR generates 4224 tokens even with a relatively small image size of 704×256). This increased token size significantly amplifies the computational cost of constructing the similarity matrix, making ToMe inefficient for 3D object detection tasks. Compared to ToMe, GPQ not only preserves the model’s performance but also achieves faster inference speed.

### v3

@@ Abstract
Query-based models are extensively used in 3D object detection tasks, with a wide range of pre-trained checkpoints readily available online. However, despite their popularity, these models often require an excessive number of object queries, far surpassing the actual number of objects to detect. The redundant queries result in unnecessary computational and memory costs. In this paper, we find that not all queries contribute equally – a significant portion of queries have a much smaller impact compared to others. Based on this observation, we propose an embarrassingly simple approach called Gradually Pruning Queries (GPQ), which prunes queries incrementally based on their classification scores. A key advantage of GPQ is that it requires no additional learnable parameters. It is straightforward to implement in any query-based method, as it can be seamlessly integrated as a fine-tuning step using an existing checkpoint after training. With GPQ, users can easily generate multiple models with fewer queries, starting from a checkpoint with an excessive number of queries. Experiments on various advanced 3D detectors show that GPQ effectively reduces redundant queries while maintaining performance. Using our method, model inference on desktop GPUs can be accelerated by up to 1.35x. Moreover, after deployment on edge devices, it achieves up to a 67.86% reduction in FLOPs and a 65.16% decrease in inference time. The code will be available at https://github.com/iseri27/Gpq.
@@ 1 Introduction
Figure 2: Selection frequency of queries (sorted in ascending order) during inference for different methods. Among these methods, PETR, PETRv2, FocalPETR, and StreamPETR are 3D object detection methods, and the other two are 2D object detection methods. As illustrated, the selection frequency of queries is imbalanced across both 2D and 3D query-based methods. In PETR, PETRv2, and FocalPETR, there are even queries that were never selected as final results.
@@ 1.2 Pruning Redundant Queries
    We conducted extensive experiments on various query-based detectors to evaluate the effectiveness of our proposed method. The results indicate that, on desktop GPUs, GPQ achieves inference acceleration of up to 1.35× and it achieves at most a 67.87% reduction in FLOPs with a 65.16% decrease in inference time after deployment on edge devices.
@@ 3.1 Revisiting Query-based Methods
$\text{Attn}=\text{Softmax}$ $\left(\cfrac{(QW^{Q})(KW^{K})^{\mathrm{T}}}{\sqrt{D_{v}}}\right)(VW^{V})W^{O}$ (1)
@@
$Q_{l}$ $\leftarrow\text{Self-Attention}_{l}(Q_{l-1})$ (2)
$Q_{l}$ $\leftarrow\text{Cross-Attention}_{l}(Q_{l},F_{I},F_{I})$
@@ 4.1 Experimental Setup
RayDN
(ECCV 2024)
ResNet50 704x256 900 / - 15.3 45.93% 0.5455 0.5880 0.2666 0.5144 0.2674 0.2042
300 / - 16.8 43.31% 0.5255 0.6020 0.2728 0.5526 0.2737 0.2093
900 / 300 17.0 45.71% 0.5507 0.5923 0.2636 0.4652 0.2590 0.1989
Table 2: Pruning results across models. The column of “Queries” is of format “# Initial queries/ # Final queries”. Lines where “# Final queries” remain blank are baselines used to compare. All baselines and checkpoints are trained for 24 epochs, with the pruning process completed within the first 6 epochs when loading a checkpoint. FPS is measured on a single RTX3090.
4.1.1 Dataset and Detectors
We conduct our experiments on the nuScenes dataset [2], which consists of over 23,000 samples. Each sample includes images from six surrounding cameras, covering the full 360 field of view. The 3D detection task involves 10 object classes (car, truck, bus, trailer, construction vehicle, pedestrian, motorcycle, bicycle, barrier, and traffic cone), including both static and dynamic objects.
To validate the effectiveness and efficiency of our proposed method, we perform experiments on five advanced detectors: DETR [49], PETR [33], PETRv2 [34], FocalPETR [47], and StreamPETR [48]. Notably, PETRv2 employs VovNet [22] as its image backbone, while the others utilize ResNet50 [15] as their image backbone.
4.1.2 Evaluation Metrics
In the field of 3D object detection, the primary performance evaluation metrics are mean Average Precision (mAP). NuScenes also provides the nuScenes Detection Score (NDS), which is derived from mAP, along with several other error metrics: mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), mean Average Velocity Error (mAVE), and mean Average Attribute Error (mAAE). Mathematically, $\text{NDS}=\cfrac{1}{10}\left[5\text{mAP}+\sum\limits_{\text{mTP}}\left(1-\min(1,\text{mTP})\right)\right]$, where $\text{mTP}\in\{\text{mATE},\text{mASE},\text{mAOE},\text{mAVE},\text{mAAE}\}$.
4.1.3 Implementation Details
We use AdamW [36] optimizer with a weight decay of $1.0\times 10^{-2}$. When training from scratch, we use a base learning rate of $2\times 10^{-4}$ with a batch size of $8$, and apply a cosine annealing policy [37] for learning rate decay. When fine-tuning an existing checkpoint, the learning rate is set to $1.0\times 10^{-4}$. The point cloud range is configured to $61.2\mathrm{m},61.2\mathrm{m}]$ along the $X$ and $Y$ axes, and $10.0\mathrm{m},10.0\mathrm{m}]$ along the $Z$ axis for all methods.
@@ 4.2 Main Results
4.2.1 Performance: Before vs. After Pruning
@@
One of the key objectives of our pruning method is to maintain the performance of the original models. As demonstrated in Table 2, our approach successfully preserves, and in some cases enhances, the performance of various detectors, even when pruning a checkpoint with a large number of queries. GPQ can also accelerate model inference on desktop GPUs. For example, in the case of PETR, pruning from 900 to 150 queries results in an mAP of 30.52%, which is 2.15 points higher than training from scratch with 150 queries (28.37%), and its speed increases from 6.9 fps to 9.3, which is 1.35x faster. Remarkably, pruning from 900 to 300 queries achieves an mAP of 32.85%, outperforming the result of training from scratch with 900 queries (31.74%). Similarly, for both FocalPETR and StreamPETR, pruning from 900 to 300 queries yields optimal performance, surpassing the results obtained from training with 900 queries from scratch. For PETRv2 and RayDN[30], while pruning from 900 to 300 queries results in a slightly lower mAP than training from scratch with 900 queries, it still exceeds the performance of training from scratch with 300 queries.
@@
Pruning r Queries Speed(fps) mAP $\uparrow$ NDS $\uparrow$ mATE $\downarrow$ mASE $\downarrow$ mAOE $\downarrow$ mAVE $\downarrow$ mAAE $\downarrow$
- - 900/- 16.1 37.83% 0.4734 0.6961 0.2822 0.6846 0.2856 0.2084
ToMe (ICLR 2023) 2112 900/- 16.0 31.69% 0.4325 0.7546 0.2907 0.6893 0.3170 0.2210
GBC (ICCV 2025) 2000 900/- 19.3 37.93% 0.4817 0.6787 0.2758 0.6390 0.2844 0.2016
GPQ - 900/300 18.7 39.42% 0.4941 0.6766 0.2711 0.5799 0.2780 0.2136
@@
Table 4: Comparison with ToMe and GBC. Results indicate that ToMe performs poorly on 3D object detection, while our GPQ effectively maintains the model’s performance and accelerates inference. $r$ indicates keys pruned by GBC or ToMe.
Inference speed is crucial for deploying models on edge devices. To verify whether pruning queries can indeed improve speed, we deploy models with GPQ on the Jetson Nano B01 to measure the model’s running time. As shown in Table 3, after pruning from 900 to 300 queries, the FLOPs of StreamPETR are reduced by 28%, and the running time decreases by 39.74%. Pruning further from 900 to 150 queries results in a 47.96% reduction in running time, making the model $1.92\times$ faster.
Since our method does not modify the image backbones, we remove the backbone modules to better illustrate the effects of our approach. Specifically, we run only the transformer decoder using randomly generated dummy inputs. For all the evaluated models, pruning from 900 queries to 300 results in saving more than half of the running time for transformer-related modules. Additionally, compared to the model with 900 queries, StreamPETR with 150 pruned queries achieves a 65.16% reduction in running time. The pruning method also proves effective for FocalPETR and PETR, as shown in Table 3.
4.2.3 Comparison
@@
We applied GBC [54] and ToMe [1] to StreamPETR and compared their performance with GPQ in Table 4. Surprisingly, ToMe slows down the model instead of accelerating it. This is likely due to the overhead of computing the similarity matrix. Unlike image classification tasks, which typically handle only a few hundred tokens, 3D object detection methods often generate a much larger number of tokens (e.g., StreamPETR generates 4224 tokens even with a relatively small image size of 704×256). This increased token size significantly amplifies the computational cost of constructing the similarity matrix, making ToMe inefficient for 3D object detection tasks. Different from ToMe, GBC can achieve a faster speed. However, compared to GPQ, it will lead to a performance drop. Compared to ToMe and GBC, GPQ both achieve better performance and faster speed.